OpenStudy (anonymous):

A particular strain of bacteria triples in population every 45 minutes. Assuming you start with a 50 bacteria in a petri dish, how many bacteria will there be in 4.5 hours? A. 33,960 B. 7015 C. 36,450 D. 12,150

OpenStudy (anonymous):

This can be modeled as an equation. so start off with y=mx+b

OpenStudy (anonymous):

Your Y is what we are trying to identify so it will still be sundown until later

OpenStudy (anonymous):

unknown. Sorry

OpenStudy (dangerousjesse):

Whoa, it's definitely not unknown!

OpenStudy (anonymous):

no i mean the ansewr is found later on.

OpenStudy (dangerousjesse):

Initial population times the growth rate to the power of (Time for final answer/ time for r)=total

OpenStudy (anonymous):

so r u saying that its 4.5*45+50=x

OpenStudy (dangerousjesse):

So \(\huge 5-\times 3^{4.5~hours~ or~270~minutes\div 45}\)

OpenStudy (dangerousjesse):

Ignore the "-"

OpenStudy (anonymous):

couldnt have done better my self:))))

OpenStudy (dangerousjesse):

So \(3^{270\div 45}\) is...?

OpenStudy (dangerousjesse):

@mada69 All you :)

OpenStudy (anonymous):

my calculator gave me an error.. give me a moment

OpenStudy (nincompoop):

this is an exponential growth and if you want to put it in linear form, you will have to do some algebraic manipulation

OpenStudy (anonymous):

i got 729

OpenStudy (nincompoop):

that does not look right

OpenStudy (dangerousjesse):

Only divide, and it'll fit in. Leave the three alone :P

OpenStudy (anonymous):

so i divide the 270 and 6?

OpenStudy (dangerousjesse):

270 and 45

OpenStudy (anonymous):

cx hehe oops my mistake i meant to say 270/45 cx i got 6

OpenStudy (nincompoop):

it is a first-order \[\frac{dN}{dt} = kN\] N is the number of microbes t is the time k is the constant integrate

OpenStudy (dangerousjesse):

There we go :)

OpenStudy (dangerousjesse):

Now solve the remaining equation (\(5\times3^6\) and bump a zero on there.

OpenStudy (anonymous):

wait i bump in the zero with the 6?

OpenStudy (dangerousjesse):

The answer was C :P \(5\times 3^6=3,645\) Add a zero \(36,450\)